OpenStudy (anonymous):

Write an equation of the line in slope intercept form, given a slope of 3 passing through (1,-9).

OpenStudy (skullpatrol):

What have you tried?

OpenStudy (anonymous):

What do you mean.

OpenStudy (skullpatrol):

What have you tried to answer the question for yourself?

OpenStudy (hba):

Well the formula for slope intercept form is, \[\huge\ (y-y_1)=m(x-x_1)\]

OpenStudy (hba):

Now trying plugging in the slope(m)=4 and (x1,y1)=(1,-9)

OpenStudy (hba):

@barnoldwrl Can you do it ?

OpenStudy (skullpatrol):

@hba please don't provide any formulas until we find out what she knows :)

OpenStudy (hba):

Well, providing formulas is just a kick-off to help users. Everyone has their way of teaching. Thankyou.

OpenStudy (anonymous):

I knew the formula thank you @hba...

OpenStudy (hba):

Well can you plug in the points @barnoldwrl

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

having trouble figuring it out though

OpenStudy (skullpatrol):

Do you understand the "y-intercept form" y=mx +b?

OpenStudy (skullpatrol):

Since this is what you are asked for in the question: "Write an equation of the line in slope intercept form"?

OpenStudy (skullpatrol):

Note: slope intercept form = "y-intercept form" y=mx +b

OpenStudy (skullpatrol):

Are you still there?

OpenStudy (skullpatrol):